OpenStudy (anonymous):

Express vector U=<-2,2> as sum of vectors p +n. Where p is parallel to v and n orthogonal to v. Vector V = <2,1>

OpenStudy (anonymous):

to determine the vector \(p\) parallel to \(v\) take their vector projection:$$\operatorname{proj}_v u=\frac{u\cdot v}{v\cdot v}v=\dots$$

OpenStudy (anonymous):

observe:

OpenStudy (anonymous):